OpenStudy (anonymous):

can someone please help me? Find the sum of the geometric sequence. 1/3 , 2/3 , 4/3, 8/3 , 16/3

OpenStudy (anonymous):

I think it's 31/3....

OpenStudy (anonymous):

the q=2 you just apply the sum rule

OpenStudy (anonymous):

21 isn't even an answer...

OpenStudy (loser66):

if you don't want apply the formula, just do : \[\dfrac{1}{3} +\dfrac{2}{3}+\dfrac{4}{3}+\dfrac{8}{3}+\dfrac{16}{3}\] they have the same denominator, wow.... easy.

OpenStudy (anonymous):

so it was 31/3! :D

OpenStudy (anonymous):

sorry,i confused this with something else

OpenStudy (loser66):

yup

OpenStudy (anonymous):

Thank you, again! c: And that's fine Wonderer16.